### v1

# SupSup Ref

Abstract
A long-standing goal of 3D human reconstruction is to create lifelike and fully detailed 3D humans from single images. The main challenge lies in inferring unknown human shapes, clothing, and texture information in areas not visible in the images. To address this, we propose SiTH, a novel pipeline that uniquely integrates an image-conditioned diffusion model into a 3D mesh reconstruction workflow. At the core of our method lies the decomposition of the ill-posed single-view reconstruction problem into hallucination and reconstruction subproblems. For the former, we employ a powerful generative diffusion model to hallucinate back appearances from the input images. For the latter, we leverage skinned body meshes as guidance to recover full-body texture meshes from the input and back-view images. Our designs enable training of the pipeline with only about 500 3D human scans while maintaining its generality and robustness. Extensive experiments and user studies on two 3D reconstruction benchmarks demonstrated the efficacy of our method in generating realistic, fully textured 3D humans from a diverse range of unseen images.
Figure 1: Single-view textured human reconstruction. We propose SiTH, a novel approach for creating high-quality and fully textured 3D human meshes from single images. We first hallucinate back-view appearances through an image-conditioned diffusion model, followed by the reconstruction of full-body textured meshes using both the front and back-view images. Our pipeline enables the creation of lifelike and diverse 3D humans from unseen photos (left) and AI-generated images (right). Best viewed in color.
1 Introduction
With the growing popularity of 3D and virtual reality applications, there has been increasing interest in creating realistic 3D human models. In general, crafting 3D humans is labor-intensive, time-consuming, and requires collaboration from highly skilled professionals. To bring lifelike 3D humans to reality and to support both expert and amateur creators in this task, it is essential to enable users to create textured 3D humans from simple 2D images or photos.
Reconstructing a fully textured human mesh from a single-view image presents an ill-posed problem with two major challenges. Firstly, the appearance information required for generating texture in unobserved regions is missing. Secondly, 3D information for mesh reconstruction, such as depth, surface, and body pose, becomes ambiguous in a 2D image. Previous efforts [44, 60, 2] attempted to tackle these challenges in a data-driven manner, focusing on generating image-mesh pairs to train neural networks. However, these approaches struggle with images featuring unseen appearances or poses (see Fig. 5), due to limited 3D human training data. More recent studies [45, 60, 54, 55] introduced additional 3D reasoning modules to enhance robustness against unseen poses. Yet, generating realistic and full-body textures from unseen appearances still remains an unsolved problem.
To address the above challenges, we propose SiTH, a novel pipeline that uniquely integrates an image-condition diffusion model to reconstruct lifelike 3D textured humans from monocular images. At the core of our approach is the decomposition of the challenging single-view problem into two subproblems: generative back-view hallucination and mesh reconstruction. This decomposition enables us to exploit the generative capability of pretrained diffusion models to guide full-body mesh and texture reconstruction. The workflow is depicted in Fig. 1. Given a front-view image, the first stage involves hallucinating a perceptually consistent back-view image using image-conditioned diffusion. The second stage reconstructs full-body mesh and texture, utilizing both the front and back-view images as guidance.
More specifically, we employ the generative capabilities of pretrained diffusion models (e.g. Stable Diffusion [41]) to infer unobserved back-view appearances for full-body 3D reconstruction. The primary challenge in ensuring the realism of 3D meshes lies in generating images that depict spatially aligned body shapes and perceptually consistent appearances with the input images. While diffusion models demonstrate impressive generative abilities with text conditioning, they are limited in producing desired back-view images using the frontal images as image conditions. To overcome this, we adapt the network architecture to enable conditioning on frontal images and introduce additional trainable components following ControlNet [58] to provide pose and mask control. To fully tailor this model to our task while retaining its original generative power, we apply adapter-based fine-tuning [15, 58] using multi-view images rendered from 3D human scans. Complementing this generative model, we develop a mesh reconstruction module to recover full-body textured mesh from front and back-view images. We follow prior work in handling 3D ambiguity through normal [45] and skinned body [54, 60] guidance. Compared to existing end-to-end methods, our two-stage pipeline can recover full-body textured meshes, including back-view details, and demonstrates robustness to unseen images (Fig. 1). It is noteworthy that the models for both subproblems are trained using the same public THuman2.0 [56] dataset, which consists of about 500 scans.
Furthermore, we create a single-view human reconstruction benchmark using a high-quality 3D human dataset, CustomHumans [14]. In this benchmark, we conducted detailed evaluations, ablations, and user studies, to compare our method with state-of-the-art approaches. Additionally, we demonstrated an application combining text-guided diffusion models, showing our potential for supporting 3D human creation. Our contributions are three-fold:
  • We propose SiTH, a single-view human reconstruction pipeline that generates lifelike and fully textured 3D humans using image-condition diffusion models.
  • By decomposing the complex, ill-posed problem, SiTH is robust to unseen images and can be effectively trained using publicly available 3D human scans.
  • SiTH can be seamlessly integrated with powerful text-to-image diffusion models for 3D human creation.
Figure 2: Method overview. SiTH is a two-stage pipeline composed of back-view hallucination and mesh reconstruction. The back-view hallucination module samples perceptually consistent back-view images through an iterative denoising process conditioned on the input image, UV map, and silhouette mask (Sec. 3.1). Based on the input and generated back-view images, the mesh reconstruction module recovers a full-body mesh and textures leveraging a skinned body prior as guidance (Sec. 3.2). Note that both modules in the pipeline can be trained with the same public 3D human dataset and generalize unseen images.
2 Related Work
Single-view human mesh reconstruction.
We review methods that recover 3D human shapes, garments, and textures. One line of work utilizes neural fields [53] to represent clothed 3D humans. As a seminal work, Saito et al. [44] first proposed a data-driven method with pixel-aligned features. Its follow-up work PIFuHD [45] further improved this framework with high-res normal guidance. Follow-up approaches extend this framework with additional human body priors. For instance, PaMIR [60] and ICON [54] utilize skinned body models [28, 35] to guide 3D reconstruction. ARCH [16], ARCH++ [13], and CAR [25] transform global coordinates into the canonical coordinates to allow for reposing. PHOHRUM [2] and S3F [7] further disentangle shading and albedo to enable relighting. Another line of work replaces the neural representations with conventional Poisson surface reconstruction [20, 19]. ECON [55] and 2K2K [10] train normal and depth predictors to generate front and back 3D point clouds. The human mesh is obtained by combining these point clouds with body priors and 3D heuristics. However, the above models cannot produce realistic full-body texture and geometry in the unobserved regions. Our pipeline addresses this problem by incorporating a generative diffusion model into single-view human reconstruction.
3D generation with 2D diffusion models.
Diffusion models [41, 43, 40, 36] trained with large collections of images have demonstrated unprecedented capability in creating 3D objects with text prompts. Most prior work [37, 51, 26, 32, 11, 6] followed an optimization workflow to update 3D representations (e.g. NeRF [33], SDF tetrahedron [46]) via neural rendering [50] and a score distillation sampling (SDS) [37] loss. While some methods [57, 5, 17] applied this workflow to human bodies, only common subjects can be generated due to the ambiguity of text-conditioning. More recent work [27, 38, 48, 47] also tried to extend this workflow with more accurate image-conditioning. However, 3D humans require geometric details on clothing and facial expressions. In our experiments, we show that these methods struggle to recover such details and take a long optimization time. Most related to our work is Chupa [21], which also decomposes the pipeline into two stages. Note that Chupa does not generate texture and only uses texts for conditioning. Our methods address these issues by leveraging an image-conditioned diffusion model to create fully textured human meshes. More importantly, our method does not require a long optimization process.
Diffusion models adaptation.
Foundation models [8, 3, 12, 22] trained on large-scale datasets have been shown to be adaptable to various downstream tasks. Following this trend, pretrained diffusion models [41, 43, 40, 36] have become common backbones for generative modeling. For instance, they can be customized by finetuning with a small collection of images [15, 23, 42]. ControlNet [58] introduced additional trainable plugins to enable image conditioning such as body skeletons. While these strategies have been widely adopted, none of them directly fit our objective. More relevant to our task is DreamPose [18], which utilizes DensePose [9] images as conditions to repose input images. However, it cannot handle out-of-distribution images due to overfitting. Zero-1-to-3 [27] finetunes a diffusion model with multi-view images to allow for viewpoint control. However, we show that viewpoint conditioning is not sufficient for generating consistent human bodies. Our model addresses this issue by providing accurate body pose and mask conditions for back-view hallucination.
3 Methodology
Method overview.
Given an input image of a human body and estimated SMPL-X [35] parameters, SiTH produces a full-body textured mesh. This mesh not only captures the observed appearances but also recovers geometric and textural details in unseen regions, such as clothing wrinkles on the back. The pipeline is composed of two modules and is summarized in Fig. 2. In the first stage, we hallucinate unobserved appearances leveraging the generative power of an image-conditioned diffusion model (Sec. 3.1). In the second stage, we reconstruct a full-body textured mesh given the input front-view image and the generated back-view image as guidance (Sec. 3.2). Notably, both modules are efficiently trained with 500 textured human scans in THuman2.0 [56].
3.1 Back-view Hallucination
Preliminaries.
Given an input front-view image $I^{F}\in\mathbb{R}^{H\times W\times 3}$, our goal is to infer a back-view image $I^{B}\in\ \mathbb{R}^{H\times W\times 3}$ which depicts unobserved body appearances. This task is under-constrained since there are multiple possible solutions to the same input images. Taking this perspective into account, we leverage a latent diffusion model (LDM) [41] to learn a conditional distribution of back-view images given a front-view image. First, a VAE autoencoder, consisting of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, is pretrained on a corpus of 2D natural images through image reconstruction, i.e. $\tilde{I}=\mathcal{D}(\mathcal{E}(I))$. Afterwards, an LDM learns to produce a latent code $z$ within the VAE latent distribution $z=\mathcal{E}(I)$ from randomly sampled noise. To sample an image, a latent code $\tilde{z}$ is obtained by iteratively denoising Gaussian noise. The final image is reconstructed through the decoder, i.e., $\tilde{I}=\mathcal{D}(\tilde{z})$.
Image-conditioned diffusion model.
Simply applying the LDM architecture to our task is not sufficient since our goal is to learn a conditional distribution of back-view images given an input conditional image. To this end, we make several adaptations to allow for image-conditioning as shown in Fig. 3. First, we utilize the pretrained CLIP [39] image encoder to extract image features from the front-view image (i.e., $I^{F}$). These image features are used for conditioning the LDM via cross-attention transformer layers, ensuring the output image shares a consistent appearance with the input image. Second, we follow the idea of ControlNet [58] and propose to use a UV map ($I^{B}_{UV}\in\mathbb{R}^{H\times W\times 3}$) and a silhouette mask ($I^{B}_{M}\in\mathbb{R}^{H\times W}$) from the back view as additional conditions. These conditional signals provide additional information that ensures the output image has a similar body shape and pose to the conditional input image.
Learning hallucination from pretraining.
Another challenge in training an image-conditioned LDM is data. Training the model from scratch is infeasible due to the requirement of a large number of paired images rendered from 3D textured human scans. Inspired by the concept of learning from large-scale pretraining [8, 12], we build our image-conditioned LDM on top of a pretrained diffusion U-Net [41]. We follow the training strategy in [23, 58] to optimize cross-attention layers and ControlNet parameters while keeping most of the other parameters frozen (see Fig. 3). The design and training strategy of our diffusion model enables hallucinating plausible back-view images that are cosistent with the frontal inputs.
Figure 3: Training of back-view hallucination module. We employ a pretrained LDM and ControlNet architecture to enable image conditioning. To train our model, we render training pairs of conditional images $I^{F}$ and ground-truth images $I^{B}$ from 3D human scans. Given a noisy image latent $z_{t}$, the model predicts added noise $\epsilon$ given the conditional image $I^{F}$, UV map $I^{B}_{UV}$, and mask $I^{B}_{M}$ as conditions. We train the ControlNet model and cross-attention layers while keeping other parameters frozen.
Training and inference.
To generate pairwise training images from 3D human scans, we sample camera view angles and use orthographic projection to render RGBA images from 3D scans and UV maps from their SMPL-X fits. Given a pair of images rendered by a frontal and its corresponding back camera, the first image serves as the conditional input $I^{F}$ while the other one is the ground-truth image $I^{B}$. During training, the ground-truth latent code $z_{0}=\mathcal{E}(I^{B})$ is perturbed by the diffusion process in t time steps, resulting in a noisy latent $z_{t}$. The image-conditoned LDM model $\epsilon_{\theta}$ aims to predict the added noise $\epsilon$ given the noisy latent $z_{t}$, the time step $t\sim[0,1000]$, the conditional image $I^{F}$, the silhouette mask $I^{B}_{M}$, and the UV map $I^{B}_{UV}$ (See Fig. 3). The objective function for fine-tuning can be represented as:
$\mathop{\min_{\theta}}\mathbb{E}_{z\sim\mathcal{E}(I),t,\epsilon\sim\mathcal{N}(\mathbf{0,I})}\left\|\epsilon-\epsilon_{\theta}(z_{t},t,I^{F},I^{B}_{UV},I^{B}_{M})\right\|_{2}^{2}.$ (1)
At test time, we obtain $I^{B}_{UV},I^{B}_{M}$ from an off-the-shelf pose predictor [4] and segmentation model [22]. To infer a back-view image, we sample a latent $\tilde{z}_{0}$ by performing the iterative denoising process starting from a Gaussian noise $z_{T}\sim\mathcal{N}(\mathbf{0,I})$. The back-view image can obtained by:
$\tilde{I}^{B}=\mathcal{D}(\tilde{z}_{0})=\mathcal{D}(f_{\theta}(z_{T},I^{F},I^{B}_{UV},I^{B}_{M})),$ (2)
where $f_{\theta}$ is a function representing the iterative denoising process of our image-conditioned LDM (See Fig. 2 left).
Figure 4: Mesh reconstruction module. Given front and back-view images ($I^{F},I^{B}$) we predict their normal images ($N^{F},N^{B}$) through a learned normal predictor. A 3D point $\mathbf{x}$ is projected onto these images for querying pixel-aligned features ($\mathbf{f}_{d,x},\mathbf{f}_{r,x}$). To leverage human body mesh as guidance, we embed the point $\mathbf{x}$ into the local UV coordinates $\mathbf{u}_{c}$, vector $\mathbf{n_{c}}$, distance $d_{c}$, and visibility $v_{c}$. Finally, two decoders ($H_{d},H_{r}$) predict SDF and RGB values at $\mathbf{x}$ given the positional embedding and pixel-aligned features.
Figure 5: Qualitative comparison on CustomHumans. Top: Results of methods generating mesh and texture. Bottom: Results of methods generating mesh only. Note that single-view reconstruction is not possible to replicate exact back-view texture and geometry. Our method generates realistic texture and clothing wrinkles perceptually close to the real scans while other baselines only produce smooth colors and surfaces in the back regions. Best viewed in color and zoom in.
3.2 Human Mesh Reconstruction
After obtaining the back-view image, our goal is to construct a full-body human mesh and its textures using the input and back-view image as guidance. We follow the literature [44, 45] to model this task with a data-driven method. Given pairwise training data (i.e., front/back-view images and 3D scans), we learn a data-driven model that maps these images to a 3D representation (e.g., a signed distance field (SDF)). We define this mapping as below:
$\mathbf{\Phi}:\mathbb{R}^{H\times W\times 3}\times\mathbb{R}^{H\times W\times 3}\times\mathbb{R}^{3}$$\rightarrow\mathbb{R}\times\mathbb{R}^{3}$ (3)
$(I^{F},I^{B},\mathbf{x})$ $\mapsto d_{x},\mathbf{r}_{x},$
where $\mathbf{x}$ is the 3D coordinate of a query point, and $d_{x},\mathbf{r}_{x}$ denote the signed distance and RGB color value at point $\mathbf{x}$. The network components we used for learning the mapping function are depicted in Fig. 4.
Local feature querying.
To learn a generic mapping function that is robust to unseen images, it is important that the model is conditioned solely on local image information with respect to the position of $\mathbf{x}$. Therefore, we employ the idea of pixel-aligned feature querying [44, 45] and separate our model into two branches, i.e., color and geometry. Our model contains a normal predictor that converts the RGB image pair ($I^{F},I^{B}$) into normal maps ($N^{F},N^{B}$). Two image feature encoders $G_{d},G_{r}$ then extract color and geometry feature maps $(\mathbf{f}_{d},\mathbf{f}_{r})\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times D}$ from the images and normal maps respectively (for simplicity we describe the process for a single image and leave out the superscripts, but both front and back images are treated the same). Finally, we project the query point $\mathbf{x}$ onto the image coordinate (Fig. 4 red points) to retrieve the local features $(\mathbf{f}_{d,x},\mathbf{f}_{r,x})\in\mathbb{R}^{D}$:
$\mathbf{f}_{d,x}=\mathcal{B}(\mathbf{f}_{d},\pi(\mathbf{x}))=\mathcal{B}(G_{d}(N),\pi(\mathbf{x})),$ (4)
$\mathbf{f}_{r,x}=\mathcal{B}(\mathbf{f}_{r},\pi(\mathbf{x}))=\mathcal{B}(G_{r}(I),\pi(\mathbf{x})),$ (5)
where $\mathcal{B}$ is a local feature querying operation using bilinear interpolation and $\mathbf{\pi}(\cdot)$ denotes orthographic projection.
Local positional embedding with skinned body prior.
As mentioned in Sec. 1, a major difficulty in mesh reconstruction is 3D ambiguity where a model has to infer unknown depth information between the front and back images. To address this issue, we follow prior work [60, 54, 14] leveraging a skinned body mesh [35] for guiding the reconstruction task. This body mesh is regarded as an anchor that provides an approximate 3D shape of the human body.
To exploit this body prior, we devise a local positional embedding function that transforms the query point $\mathbf{x}$ into the local body mesh coordinate system. We look for the closest point $\mathbf{x}_{c}^{*}$ on the body mesh (Fig. 4 blue point), i.e.,
$\mathbf{x}_{c}^{*}=\arg\min_{\mathbf{x}_{c}}\|\mathbf{x}-\mathbf{x}_{c}\|_{2},$ (6)
where $\mathbf{x}_{c}$ are points on the skinned body mesh $\mathcal{M}$. Our positional embedding $\mathbf{p}$ constitutes four elements: a signed distance value $d_{c}$ between $\mathbf{x}_{c}^{*}$ and $\mathbf{x}$, a vector $\mathbf{n}_{c}=(\mathbf{x}-\mathbf{x}_{c}^{*})$, the UV coordinates $\mathbf{u}_{c}\in[0,1]^{2}$ of the point $\mathbf{x}_{c}^{*}$, and a visibility label $v_{c}\in\{1,-1,0\}$ that indicates whether $\mathbf{x}_{c}^{*}$ is visible in the front/back image or neither. Finally, two separate MLPs $H_{d},H_{r}$ take the positional embedding $\mathbf{p}=[d_{c},\mathbf{n}_{c},\mathbf{u}_{c},v_{c}]$ and the local texture/geometry features $(\mathbf{f}_{d,x},\mathbf{f}_{r,x})$ as inputs to predict the final SDF and RGB values at point $\mathbf{x}$:
$d_{x}=H_{d}(\mathbf{f}_{d,x}^{F},\mathbf{f}_{d,x}^{B},\mathbf{p}),$ (7)
$\mathbf{r}_{x}=H_{r}(\mathbf{f}_{r,x}^{F},\mathbf{f}_{r,x}^{B},\mathbf{p}).$ (8)
CAPE [30] CustomHumans [14]
[…]
PIFu* [44] 3.350 0.755 29.283 0.9898 0.0198 2.939 0.780 31.911 0.9675 0.0454
PIFuHD* [45] 3.185 0.755 32.157 - - 2.651 0.785 36.311 - -
PaMIR [60] 3.405 0.742 25.475 0.9881 0.0140 3.591 0.753 25.716 0.9640 0.0348
ICON [54] 3.240 0.757 25.898 - - 3.190 0.763 26.857 - -
ECON [55] 3.084 0.765 26.960 - - 3.315 0.757 27.075 - -
SiTH (Ours) 2.530 0.790 31.256 0.9926 0.0111 2.443 0.805 33.061 0.9732 0.0291
Table 1: Single-view human reconstruction benchmarks. We report Chamfer distance (CD), normal consistency (NC), and f-score between ground truth and predicted meshes. To evaluate texture reconstruction quality, we compute SSIM and LPIPS between the input and front-view texture rendering. The best and the second best methods are highlighted in bold and underlined respectively. Note that * denotes models that have been trained on commercial 3D human scans while the others are trained on the public THuman2.0 dataset.
Training and inference.
We used the same 3D dataset described in Sec. 3.1 to render training image pairs $(I^{F},I^{B})$ from the 3D textured scans. For each training scan, query points $\mathbf{x}$ are sampled within a 3-dimensional cube $[-1,1]^{3}$. For each point, we compute the ground-truth signed distance values $d$ to the scan surface, closest texture RGB values $r$, and surface normal $\mathbf{n}$. Finally, we jointly optimized the normal predictors, the image encoders, and the MLPs in both branches with the following reconstruction losses:
$\mathcal{L}_{d}=\|d-d_{x}\|_{1}+\lambda_{n}(1-\mathbf{n}\cdot\nabla_{\mathbf{x}}d_{x}),$ (9)
$\mathcal{L}_{r}=\|\mathbf{r}-\mathbf{r}_{x}\|_{1}.$ (10)
Note that $\nabla_{\mathbf{x}}$ indicates numerical finite differences at point $\mathbf{x}$ and $\lambda_{n}$ is a hyperparameter.
During inference, we use the input image $I^{F}$ and the back-view image $\tilde{I}^{B}$ obtained from Sec. 3.1 to reconstruct 3D mesh and textures. First, we align both images with the estimated body mesh $\mathcal{M}$ to ensure that image features can be properly queried around the 3D anchor. We adopt a similar strategy with [54] to optimize the scale and the offset of the body mesh with silhouette and 2D joint errors. Finally, we perform the marching cube algorithm [29] by querying SDF and RGB values within a dense voxel grid via Eq. 3 (see Fig. 2 right).
Figure 6: Qualitative comparison of back-view hallucination. We visualize back-view images generated by the baseline methods. Note that the three different images are sampled from different random seeds. Our results are perceptually close to the ground-truth image in terms of appearances and poses. Moreover, our method also preserves generative stochasticity for handling tiny wrinkle changes.
4 Experiments
4.1 Experimental Setup
Dataset.
Previous work [2, 44, 45] relied on training data from commercial datasets such as RenderPeople [1]. While these datasets offer high-quality textured meshes, they also limit reproducibility due to limited accessibility. For fair comparisons, we follow ICON [54] by training our method on the public 3D dataset THuman2.0 [56] and using the CAPE [30] dataset for evaluation. However, we observed potential biases in the evaluation due to the low-res ground-truth meshes and image rendering defects in the CAPE dataset (for a detailed discussion, please refer to Supp.). Consequently, we further create a new benchmark that evaluates the baselines on a higher-quality 3D human dataset CustomHumans [14]. In the following, we provide a summary of the datasets used in our experiments:
  • THuman2.0 [56] contains approximately 500 scans of humans wearing 150 different garments in various poses. We use these 3D scans to train our pipeline.
  • CAPE [30] contains 15 subjects in 8 types of tight outfits. The test set, provided by ICON, consists of 300 images rendered from three view angles. We use this dataset to conduct the 3D quantitative experiment (Sec. 4.2).
  • CustomHumans [14] contains more than 600 high-quality scans of 80 subjects in 120 different garments and varied poses. We randomly selected 60 subjects and rendered 240 test images from four view angles. We use this test data for all quantitative experiments, user studies, and ablation studies. (Sec. 4.2Sec. 4.4)
Evaluation protocol.
We follow the evaluation protocol in OccNet [31] and ICON [54] to compute 3D metrics Chamfer distance (CD), normal consistency (NC), and f-Score [49] on the generated meshes. To evaluate mesh texture, we report structure similarity (SSIM[52], LPIPS [59]. In user studies, 30 participants rank the meshes obtained by four different methods. We report the average ranking ranging from 1 (best) to 4 (worst).
4.2 Single-view Human Reconstruction
Benchmark evaluation.
We compared SiTH with state-of-the-art single-view human reconstruction methods, including PIFu [44], PIFuHD [45], PHORHUM [2], PaMIR [60], ICON [54], and ECON [55] on CAPE and CustomHumans. Note that PHORHUM is only used for qualitative comparison since a different camera system is used, leading to the misalignment with ground-truth meshes. We visualize the generated mesh texture and normals in Fig. 5. Existing methods result in over-smoothed texture and normals, particularly in the back. Our method not only generates photorealistic and perceptually consistent appearances in unobserved regions but also recovers underlying geometric details like clothing wrinkles.
The quantitative results are summarized in Tab. 1. It’s worth noting that PIFu and PIFuHD are trained with commercial datasets (denoted as * in Tab. 1), while the others are trained on the public dataset THuman2.0. To evaluate the methods leveraging a skinned body prior (i.e., PaMIR, ICON, ECON, and Ours), we follow the same pose alignment procedure proposed in the original paper for a fair comparison. Results in Tab. 1 show that methods using a body prior (ICON, ECON) outperform end-to-end methods (PIFu, PIFuHD) on tight clothing and challenging poses in CAPE. However, they fall short in handling diverse outfits in CustomHumans. On the other hand, end-to-end methods trained on commercial datasets show better performance in this respect. Notably, our method is robust across both benchmarks, achieving performance comparable to methods trained on high-quality commercial scans.
Figure 7: Comparison with optimization-based methods. Compared to methods that utilize diffusion models for optimization, our result is closer to the input image and contains local geometric details. Note that Chupa is not conditioned on the image, and fails to generate correct clothing from the text prompts.
Compared with optimization-based methods.
We also compared SiTH with methods that use pretrained diffusion models and a score distillation sampling loss [37] to optimize 3D meshes. In the case of Zero-1-to-3 [27], we used the input image to optimize an instant NGP [34] radiance field, and for Magic-123 [38], we provided additional text prompts to optimize an SDF tetrahedron [46]. From Fig. 7, we see that while both methods can handle full-body textures, they struggle with reasoning the underlying geometry and clothing details. It is worth noting that Zero-1-to-3 and Magic-123 require 10 minutes and 6 hours in optimization, respectively, while our method takes less than 2 minutes to run a marching cube of 512 resolution. Additionally, more similar to our method is Chupa [21], which generates front/back-view normals for mesh reconstruction. Note that Chupa is not conditioned on images and does not generate texture. Instead, we provided body poses and text prompts generated by an image-to-text interrogator [24] as their conditional inputs. From Fig. 7, it’s clear that text-conditioning is less accurate than image-conditioning, and the method struggles to produce diverse clothing styles such as coats. By contrast, our method reconstructs correct clothing geometry and texture through image conditioning.
ICON ECON PIFuHD Ours
Front Geometry 3.127 2.720 2.607 1.547
Back Geometry 3.193 2.513 3.093 1.200
Similarity 3.093 2.660 2.780 1.456
PIFu PaMIR PHOHRUM Ours
Front Texture 3.067 2.153 3.450 1.327
Back Texture 3.450 2.355 3.140 1.054
Similarity 3.307 2.192 3.416 1.093
Chupa Ours
User Preference 36.0% 64.0%
Table 2: User study results. Top: 30 users are asked to rank the quality of surface normal images from best (1) to worst (4). We report the average ranking of each method. Middle: Similar to the first task, users are asked to rank the quality of RGB textures. Bottom: We ask users to choose the mesh with a better quality.
User study.
The above metrics may not fully capture the quality of 3D meshes in terms of realism and local details. To address this, we conducted a user study to compare the texture and geometry quality among various baselines. We invited 30 users to rank the front/back-view texture and normal renderings of 3D meshes generated by four different methods. Additionally, we asked the users to assess the similarity between the input images and the generated meshes. The results (Tab. 2) support our claim that existing methods struggle to efficiently generate desirable back-view texture/geometry from single-view images. Our method, which leverages the generative capability of diffusion models, consistently outperforms each baseline. It also produces more preferred front-view textures and geometries, as evidenced by higher user rankings. We also conducted a user study with Chupa (in Tab. 2 bottom) which also indicates more users prefer the 3D meshes generated by our method.
[…]
W/o Body Mesh 2.962 0.780 30.714
W/o Hallucination 2.490 0.815 31.551
Full Pipeline 2.443 0.805 33.061
W/ GT Body Mesh 1.172 0.891 58.858
W/ GT Body and $I^{B}$ 1.059 0.914 63.356
Table 3: Ablation study on CustomHumans. We ablate the hallucination module and the skinned body mesh in our pipeline. Please refer to our discussion in Sec. 4.4.
Figure 8: Ablation study. We visualize back and side-view rendering of the reconstructed meshes. Our full model produced a correct body shape and more realistic clothing geometry.
4.3 Generative Capability
Image quality comparison.
Our hallucination module is a unique and essential component that generates spatially aligned human images to guide 3D mesh reconstruction. Given that our focus is on back-view hallucination, we compare the quality of generated images with the relevant generative methods in Fig. 6. A method closely related to ours is DreamPose [18], which conditions the model with DensePose images and fine-tunes the diffusion model with paired training data. However, their model tends to overfit the training data and struggles to generalize to unseen images, in contrast to our approach. While Zero-1-to-3 [27] can generalize to unseen images, their method faces challenges in generating consistent body poses given the same back-view camera. Moreover, we designed a naïve baseline that provides ControlNet [58] for corresponding text prompts using an image-to-text interrogator [24]. However, without proper image conditioning and fine-tuning, such a method cannot generate images that faithfully match the input appearances. Our method not only addresses these issues but also handles stochastic appearances (e.g., tiny differences in wrinkles) from different random seeds. We report 2D generative evaluation metrics in Supp.
3D reconstruction plugin.
We demonstrate that our hallucination module can be seamlessly integrated into existing single-view clothed human reconstruction pipelines. We implemented variants of ECON, and PIFuHD by providing them back normal from our generated back-view images. Please refer to our experiments in Supp.
4.4 Ablation Study
We conducted controlled experiments to validate the effectiveness of our proposed modules. As shown in Fig. 8, the skinned body mesh is a crucial component for 3D human reconstruction. Without this body mesh as a 3D anchor, the output mesh contains an incorrect body shape due to the depth ambiguity issue. Conversely, removing the hallucination module has minimal impact on 3D reconstruction metrics, though it slightly degrades normal consistency. However, the overall quality in both texture and geometry is incomparable with our full model (see Fig. 8 right). This is consistent with our findings in user studies, indicating that 3D metrics may not accurately reflect the perceptual quality of 3D meshes in terms of geometry. Finally, we tested two additional variants, leveraging ground-truth body meshes and real back-view images in our full pipeline, representing the upper bound of our method. As shown in Tab. 3 bottom, this additional information notably improves the 3D metrics. These results highlight the persistent challenges in the single-view reconstruction problem, including pose ambiguity and the stochastic nature of clothing geometry.
4.5 Applications
Figure 9: Applications. Our pipeline can be incorporated with text-to-image diffusion models for 3D human creation.
Inheriting the powerful generative capability from LDM, SiTH is robust to diverse input images, such as out-of-distribution or AI-generated images. We demonstrate a unique solution to link photo-realistic AI photos and high-fidelity 3D humans. In Fig. 9, we introduce a 3D creation workflow integrating powerful text-to-image diffusion models. Given an SMPL-X body mesh, we generate a front-view image using Stable Diffusion and ControlNet models with text prompts. SiTH then creates a full-body 3D model from the AI-generated images.
5 Conclusion
We propose an innovative pipeline designed to create fully textured 3D humans from single-view images. Our approach seamlessly integrates an image-conditioned diffusion model into the existing data-driven 3D reconstruction workflow. Leveraging the generative capabilities of the diffusion model, our method efficiently produces lifelike 3D humans from a diverse range of unseen images, trained with only 500 human scans. We expect our work will advance the application of generative AI in 3D human creation.
Acknowledgements
We thank Xu Chen for insightful discussions, Manuel Kaufmann for suggestions on writing and the title, and Christoph Gebhardt, Marcel Buehler, and Juan-Ting Lin for their writing advice.